OpenStudy (anonymous):

You invest an initial $100 in an account that has an annual interest rate of 3%, compounded quarterly. How much money will you have in the account after 20 years? Round your answer to the nearest whole number

OpenStudy (misty1212):

HI!! again

OpenStudy (anonymous):

haha hi. thank you. ur the most help

OpenStudy (misty1212):

\[100\left(1.03\right)^{20}\]should work

OpenStudy (anonymous):

Thank you so mch!

OpenStudy (mathstudent55):

Start by writing the formula for compound interest. Then write in your values. Then calculate.

OpenStudy (misty1212):

wrong i was wrong wrong wroing

OpenStudy (misty1212):

Quarterly !!

OpenStudy (anonymous):

???

OpenStudy (misty1212):

\[100\times (1+\frac{.03}{4})^{80}\] is what you need

OpenStudy (anonymous):

oh ahha thank you

OpenStudy (misty1212):

formula \[P(1+\frac{r}{n})^{nt}\] is what i am using

OpenStudy (misty1212):

P is the principle, r is the rate (as a decimal not as a percent) n is the number of compounding periods and t is time in years

OpenStudy (rational):

looks like you would have lost more than $1 with that wrong formula http://www.wolframalpha.com/input/?i=100%281%2B.03%2F4%29^%2880%29+-+100%281%2B.03%29^%2820%29

OpenStudy (mathstudent55):

\(F = P \left(1 + \dfrac{r}{n} \right) ^{nt} \) where F = future value P = present value r = annual interest rate of interest (written as a decimal) n = number of times the interest is compounded per year t = number of years

OpenStudy (mathstudent55):

\(F = 100 \left(1 + \dfrac{0.03}{4} \right) ^{4 \times 20}\)

OpenStudy (mathstudent55):

\(F = 100 (1.075) ^{80}\) \(F = 100(1.81804...)\) \(F = 181.80\)